Hydrofluoric acid has a Ka value of 7.2x10−4. What is the pH of a 0.04M solution

of hydrofluoric acid?


1
Expert's answer
2022-03-08T10:55:01-0500

"7.2×10^{-4}=\frac{[H^+][F^-]}{[HF]}"


"=\frac{(x)(x)}{(0.04-x)}"


Ignoring "x" in the denominator

"x^2=2.88\u00d710^{-5}"

"x=5.367\u00d710^{-3}"


pH"=-" log "(5.367\u00d710^{-3})"

"=2.3"